Caffeine, a stimulant found in coffee and some soft drinks, consists of 49.5% C, 5.15% H, and 28.9% N by mass; the rest is oxygen. Given the molar mass of caffeine is 195 g/mol, determine its molecular formula.


1
Expert's answer
2022-06-06T17:04:11-0400

M(C) = 12 g/mol

M(H) = 1 g/mol

M(N) = 14 g/mol

M(O) = 16 g/mol

To find the empirical formula, we divide the percentage of each element by its atomic mass. we divide the result by the smallest of the results. if whole numbers come out, this is the empirical formula. if an integer does not appear, multiply each of them by 2; 3, and so on until an integer is formed.

C– 49,5/12 = 4,125 /1,03= 4

H– 5,15/1 = 5,15 /1,03= 5

N– 28,9/14 = 2,06 /1,03= 2

O– 16,45/16 = 1,03 /1,03=1

That is, C4H5N2O = 97 g

Multiply each by 2 so that the molecular mass is 195 g/mol As a result: C8H10N4O2


Answer: C8H10N4O2